# Simulations of dielectric permittivity of water by Machine Learned Potentials with long-range Coulombic interactions

### v1

@@ I Introduction
In theory, the dielectric constant of water can be calculated based on the knowledge of dipole fluctuation in an equilibrated liquid trajectory within the linear response regime [5]. However, the computer simulations of dielectric constant in water have faced significant challenges over decades. The difficulty lies in the fact that the dipole fluctuation in water is collective in nature which is hard to converge on both time and space scales. In the time domain, a converged dipole fluctuation in water needs to include the oscillation of electric dipoles on water molecules in response to an electric field change, which is characterized by the dielectric relaxation time or referred to as Debye relaxation time. In experiments, the dielectric relaxation time is estimated to be as large as 8-9 picoseconds (psps) because of the underlying Hydrogen(H)-bond network. To sufficiently sample the dielectric relaxations in the calculated dipole fluctuations, the liquid trajectories are required to be simulated at a time scale of tens of nanoseconds (nsns). In the space domain, the theoretical determination of dielectric constant requires to converge the correlated dipole fluctuations on the underlying H-bond network as well. Because of the directional H-bonding, a water dipole is highly correlated with the dipole on a neighboring water molecule. On a near tetrahedral H-bond network of liquid water, this dipole-dipole correlation decays slowly as the distance between the pair of molecules increases, and a pair of dipoles remain correlated at the scale of nanometers which is characterized by the Kirkwood correlation length. As such, an accurate modeling of dielectric constant demands the modeling of the collective fluctuations of correlated dipoles in real space. In this regard, the liquid water is required to be modeled by large supercell whose size should be comparable to its correlation length.
@@ II.2 Deep Neural Network (DNN) Models
To explore the potential energy surface, PI-DPMD simulations were performed in the isobaric-isothermal (NPTNPT) ensemble at 1 bar and 330K or 300 K with 64 water molecules using the i-PI code [36] interfaced with the DeePMD-kit package [34]. At each time step, all four independently trained DP models were used to evaluate the atomic forces, while only one randomly selected model was employed to propagate the trajectory. As the simulations progress, differences among the four models in their predictions of energy $E$ and forces $\bm{F}_{i}$ inevitably emerge. To assess model consistency and convergence, we monitored the maximum standard deviation of the predicted forces, defined as $\zeta=\max_{i}\sqrt{\langle\left|\bm{F}_{i}-\bar{\bm{F}}_{i}\right|^{2}\rangle}$, where $\bar{\bm{F}_{i}}=\langle\bm{F}_{i}\rangle$ denotes the average force on atom $i$ over the four models. The DP models were considered converged when the fraction of configurations with $\zeta>0.2$ eV/Å dropped below $0.005\%$ of the total PI-DPMD configurations. If this threshold was not met, 50-400 high-discrepancy configurations ($\zeta>0.2$ eV/Å) were selected, and their energies, atomic forces, and WC coordinates were computed using the aforementioned DFT protocol. These new data were added to the training set, and the model training cycle was repeated until convergence was achieved.
@@ II.3 Classical DPLR-MD Simulations
The SCAN0-based DPLR model was applied to conduct DPLR-MD simulations in the NPTNPT ensemble at 1 bar and 330 K with a simulation cell containing 4096 water molecules unless otherwise specified. The DPLR-MD simulations were performed by using LAMMPS [37] interfaced with the DeePMD-kit package [34] for a total duration of 20 ns, with the initial 100 ps excluded from analysis to allow for thermal and structural equilibration. The DW model was subsequently used to compute the dielectric properties of the simulated systems. Specifically, it was applied to infer the instantaneous dipole moments of individual water molecules from the atomic configurations generated during the DPLR-MD simulations. These molecular dipoles were then used to evaluate macroscopic polarization and its fluctuations, which form the basis for calculating the static dielectric constant and related electrostatic response functions. This approach enables an accurate and efficient characterization of dielectric behavior, leveraging the electronic structure information encoded in the DW model.
@@ III Dipole Fluctuation Method
For the liquid water modeled by molecular dynamics in periodic simulation cells containing NN water molecules, the static dielectric constant $\varepsilon$ of water can be calculated via the fluctuation of the overall electric dipole moment $\mathbf{M}$ as presented in Eq. 2 based on linear response theory. In Eq. 2, $V$, $k_{\mathrm{B}}$, and $T$ are the volume of the simulation cell, Boltzmann constant, and temperature, respectively. In analogy to the experimental measurements, a theoretical simulation of polar liquid demands the specification of its electric boundary condition that being applied on water. In liquid water, the Hamiltonian, thermodynamic enthalpy energy, and predicted integral of motion are all crucially dependent on its electric boundary conditions. According to the Onsager reaction field theory [22], the dependence of the dielectric constant of an isotropic liquid on its electric boundary condition can be simply determined by one parameter $\varepsilon^{\prime}$ in Eq. 2.
@@ III.1 Dipole fluctuation under metallic boundary condition
Within the framework of the dipole fluctuation approach, the metallic electric boundary condition has been the most applied one to evaluate the dielectric constant in liquid water because most classical MD and AIMD simulations adopt Ewald summation technique in computing the electrostatic potential energies. In Ewald summation, the applied electric field on a system is constrained to zero by the vanishing electric potential drop across the macroscopic sample, which is equivalent to applying the metallic boundary condition. Under the metallic boundary condition, the dipole fluctuation of water is transverse in nature. Therefore, a converged simulation of transverse dipole fluctuation in water is essential for an accurate prediction of its static dielectric constant. As shown in Fig. 2 (a), we present the electric dipole $\mathbf{M}$ as a function of simulation time along one of the orthogonal directions in the DPMD simulations. Clearly, it can be seen in Fig. 2 (a) that $\mathbf{M}$ has a mean value $\mathbf{M}=0$, which is expected since water is a paraelectric material with vanishing net polarization at room temperature. Nevertheless, $\mathbf{M}$ undergoes a large fluctuation and a significant second moment $\left\langle\mathbf{M}^{2}\right\rangle$ giving rise to its large static dielectric constant. However, $\left\langle\mathbf{M}^{2}\right\rangle$ converges rather slowly with simulation time as shown in Fig. 2 (c), which is due to the slow relaxation of polarization. According to Debye’s theory of dielectric relaxation, the polarization fluctuation decays exponentially over time whose characteristic time $\tau^{\epsilon^{\prime}=\infty}_{\mathbf{M}}$ can be determined from the autocorrelation $C_{\mathbf{M}}\left(t\right)$ of the electric dipole in the system defined as $C_{\mathbf{M}}\left(t\right)=\frac{\left\langle\mathbf{M}\left(t\right)\cdot\mathbf{M}\left(0\right)\right\rangle}{\left\langle\mathbf{M}\left(0\right)^{2}\right\rangle}$. In our simulation, a relaxation time $\tau^{\epsilon^{\prime}=\infty}_{\mathbf{M}}=8.1$ psps has been extracted from our DPMD simulation, which is close to the experimentally reported Debye relaxation time of 8.28 psps at room temperature [40]. In the MD computer simulations, the simulation time $t$ should be long enough to sample sufficient cycles of polarization relaxation in water for a well-converged dipole fluctuation $\left\langle\mathbf{M}^{2}\right\rangle$. The formulas of the relative error $I_{\varepsilon-1}$ in the predicted dielectric constant are shown in Eq. 3 in which the accuracy is inversely proportional to $\frac{t}{\tau_{\mathbf{M}}}$ as expected. Because of the slow Debye relaxation, it has been widely noticed that simulation time at the scale of tens of nanoseconds is required to have a converged prediction of the static dielectric constant of water. In our DPMD simulation over 20 nsns , the dielectric constant of water is determined to be $\varepsilon=102$ with a small relative error of $3\%$ according to Eq. 3 [22].
@@ III.2 Dipole fluctuation under insulating boundary condition
The insulating electric boundary condition is equivalent to an open-circuit capacitor setup in experiments as schematically shown in Fig. 1 (b), in which no free charge exists on the metal plates. Because of the vanishing free charge on the capacitor, a constrained zero electric displacement field is imposed on the macroscopic sample of water [41]. Under the insulating boundary condition, the dipole fluctuation is longitudinal in nature. In comparison with the transverse dipole fluctuation under metallic boundary condition, the longitudinal dipole fluctuation is largely suppressed as shown in Fig. 2 (a). The mechanism is similar to that behind the experimentally observed higher phonon energies of longitudinal optical phonons compared to the transverse optical in ionic solid materials [42]. In addition to the suppressed dipole fluctuation, the polarization relaxation time $\tau^{\epsilon^{\prime}=0}_{\mathbf{M}}$ is greatly reduced as well. In a large band gap insulator such as neat water, no excess free charge $\rho_{\text{free}}$ exists, therefore $\nabla\cdot\mathbf{D}=\rho_{\text{free}}=0$ resulting in a uniform electric displacement field $\mathbf{D}$ throughout the system. The application of insulating electric boundary condition in DPMD simulation enables us to treat the long-range Coulombic interaction by the uniform parameter $\mathbf{D}$ throughout the neat water. Under this treatment, the dipole fluctuation in water becomes nearsighted and less correlated in real space, which results in a largely suppressed polarization relaxation time $\tau^{\epsilon^{\prime}=0}_{\mathbf{M}}$. As shown in Fig. 2 (b), the polarization relaxation time $\tau^{\epsilon^{\prime}=0}_{\mathbf{M}}=0.3$ ps is obtained in our computer simulations under the insulating electric boundary condition, and this value is in close agreement with the recent result reported by Sprik et al. [43]. Under insulating electric boundary condition, the fast polarization relaxation time $\tau^{\epsilon^{\prime}=0}_{\mathbf{M}}$ also helps the quicker convergence of $\left\langle\mathbf{M}^{2}\right\rangle$ as compared to that under metallic boundary conduction as seen in Fig. 2 (c). Because of the faster polarization relaxation time $\tau_{\mathbf{M}}^{\varepsilon^{\prime}}$ in insulating boundary condition, one expects that the predicted static dielectric constant of water should converge much faster than that in metallic boundary condition. However, our DPMD simulation suggests the opposite. As can be clearly seen in Fig. 2 (d), the predicted dielectric at the time scale of 20 nsns is barely converged with significant error bars. This seemingly surprising result is due to the fact that the prediction of $\varepsilon$ under insulating boundary condition demands the inversion of a small numerical number in Eq. 2 as $\varepsilon=\frac{1}{1-\frac{4\pi\left\langle\mathbf{M}^{2}\right\rangle}{3Vk_{\mathrm{B}}T}}$ which poses an additional challenge on the accuracy of predicted $\varepsilon$. The above effect can be more easily seen in the analytic formula of the relative errors in Eq. 2. The relative error of the insulating boundary condition has an extra term of $\varepsilon$ reflecting the loss of accuracy by inverting a small number. Based on Eq. 3, it can be estimated that a DMPD trajectory over 2.2 microseconds ($\mu s$) is required in order to achieve a similar accuracy as reported for a DPMD simulation performed under the metallic boundary condition within 20 ns.
@@ IV Dipole Correlation Method
In the above, the dipolar pair correlation function $c_{m}^{\varepsilon^{\prime}}\left(r\right)$, as defined in Eq. 6, is an ensemble function of distance $r$ normalized by the molecule number density $\rho$ and total number of molecules NN in the simulation cell, which accounts for the angular correlation between the unit vector of two dipole moments $\hat{\bm{\mu}}_{i}$ and $\hat{\bm{\mu}}_{j}$ on two distinct water molecules located at distances of $\bm{r}_{i}$ and $\bm{r}_{j}$ from the reference molecule at the origin point, respectively.
@@ IV.2 Long-Range Dipole Correlations in the H-Bond Network
The behavior of the dipolar correlation function $c_{m}^{\varepsilon^{\prime}}\left(r\rightarrow\infty\right)$ is a direct consequence of the long-range Coulomb forces. Therefore, an explicit treatment of the long-range electrostatic forces in the underlying deep neural network is essential to predict the correct asymptotic behavior. As shown in Fig. 5 (a) and (b), we compare the dipolar correlation functions $c_{m}^{\varepsilon^{\prime}=\infty}\left(r\right)$ generated by molecular simulations using the long-range Coulomb interaction inclusive DPLR model and the standard short-range DPSR model under the metallic boundary condition. In both DPLR and DPSR based molecular dynamics simulations, the liquid water structure is modeled by various sizes of simulation boxes of lengths $L$ that contain NN number of water molecules. Clearly, it can be seen that the DPLR model correctly predicts the constant of $c_{m}^{\varepsilon^{\prime}=\infty}\left(r\right)$ as $r\rightarrow\infty$. Moreover, the value of $c_{m}^{\varepsilon^{\prime}=\infty}\left(r\rightarrow\infty\right)$ decays with the volume of the simulation box, which scales as $\sim 1/V=1/L^{3}$ as expected, shown in the insert of Fig. 5 (a). In sharp contrast, $c_{m}^{\varepsilon^{\prime}=\infty}\left(r\rightarrow\infty\right)$ predicted by the DPSR model is qualitatively incorrect, yielding an exponential decay as $r\rightarrow\infty$ for all simulations, as shown in Fig. 5 (b). Indeed, in the absence of long-range Coulombic forces, the dipole correlation function, as an electronic property, exhibits exponential decay with distance from a local perturbation, in accordance with Walter Kohn’s nearsightedness principle [44].

### v2

@@ I Introduction
In theory, the dielectric constant of water can be calculated based on the knowledge of dipole fluctuation in an equilibrated liquid trajectory within the linear response regime [5]. However, the computer simulations of dielectric constant in water have faced significant challenges over decades. The difficulty lies in the fact that the dipole fluctuation in water is collective in nature which is hard to converge on both time and space scales. In the time domain, a converged dipole fluctuation in water needs to include the oscillation of electric dipoles on water molecules in response to an electric field change, which is characterized by the dielectric relaxation time or referred to as Debye relaxation time. In experiments, the dielectric relaxation time is estimated to be as large as 8-9 picoseconds ($ps$) because of the underlying Hydrogen(H)-bond network. To sufficiently sample the dielectric relaxations in the calculated dipole fluctuations, the liquid trajectories are required to be simulated at a time scale of tens of nanoseconds ($ns$). In the space domain, the theoretical determination of dielectric constant requires to converge the correlated dipole fluctuations on the underlying H-bond network as well. Because of the directional H-bonding, a water dipole is highly correlated with the dipole on a neighboring water molecule. On a near tetrahedral H-bond network of liquid water, this dipole-dipole correlation decays slowly as the distance between the pair of molecules increases, and a pair of dipoles remain correlated at the scale of nanometers which is characterized by the Kirkwood correlation length. As such, an accurate modeling of dielectric constant demands the modeling of the collective fluctuations of correlated dipoles in real space. In this regard, the liquid water is required to be modeled by large supercell whose size should be comparable to its correlation length.
@@ II.2 Deep Neural Network (DNN) Models
To explore the potential energy surface, PI-DPMD simulations were performed in the isobaric-isothermal ($NPT$) ensemble at 1 bar and 330K or 300 K with 64 water molecules using the i-PI code [36] interfaced with the DeePMD-kit package [34]. At each time step, all four independently trained DP models were used to evaluate the atomic forces, while only one randomly selected model was employed to propagate the trajectory. As the simulations progress, differences among the four models in their predictions of energy $E$ and forces $\bm{F}_{i}$ inevitably emerge. To assess model consistency and convergence, we monitored the maximum standard deviation of the predicted forces, defined as $\zeta=\max_{i}\sqrt{\langle\left|\bm{F}_{i}-\bar{\bm{F}}_{i}\right|^{2}\rangle}$, where $\bar{\bm{F}_{i}}=\langle\bm{F}_{i}\rangle$ denotes the average force on atom $i$ over the four models. The DP models were considered converged when the fraction of configurations with $\zeta>0.2$ eV/Å dropped below $0.005\%$ of the total PI-DPMD configurations. If this threshold was not met, 50-400 high-discrepancy configurations ($\zeta>0.2$ eV/Å) were selected, and their energies, atomic forces, and WC coordinates were computed using the aforementioned DFT protocol. These new data were added to the training set, and the model training cycle was repeated until convergence was achieved.
@@ II.3 Classical DPLR-MD Simulations
The SCAN0-based DPLR model was applied to conduct DPLR-MD simulations in the $NPT$ ensemble at 1 bar and 330 K with a simulation cell containing 4096 water molecules unless otherwise specified. The DPLR-MD simulations were performed by using LAMMPS [37] interfaced with the DeePMD-kit package [34] for a total duration of 20 ns, with the initial 100 ps excluded from analysis to allow for thermal and structural equilibration. The DW model was subsequently used to compute the dielectric properties of the simulated systems. Specifically, it was applied to infer the instantaneous dipole moments of individual water molecules from the atomic configurations generated during the DPLR-MD simulations. These molecular dipoles were then used to evaluate macroscopic polarization and its fluctuations, which form the basis for calculating the static dielectric constant and related electrostatic response functions. This approach enables an accurate and efficient characterization of dielectric behavior, leveraging the electronic structure information encoded in the DW model.
@@ III Dipole Fluctuation Method
For the liquid water modeled by molecular dynamics in periodic simulation cells containing $N$ water molecules, the static dielectric constant $\varepsilon$ of water can be calculated via the fluctuation of the overall electric dipole moment $\mathbf{M}$ as presented in Eq. 2 based on linear response theory. In Eq. 2, $V$, $k_{\mathrm{B}}$, and $T$ are the volume of the simulation cell, Boltzmann constant, and temperature, respectively. In analogy to the experimental measurements, a theoretical simulation of polar liquid demands the specification of its electric boundary condition that being applied on water. In liquid water, the Hamiltonian, thermodynamic enthalpy energy, and predicted integral of motion are all crucially dependent on its electric boundary conditions. According to the Onsager reaction field theory [22], the dependence of the dielectric constant of an isotropic liquid on its electric boundary condition can be simply determined by one parameter $\varepsilon^{\prime}$ in Eq. 2.
@@ III.1 Dipole fluctuation under metallic boundary condition
Within the framework of the dipole fluctuation approach, the metallic electric boundary condition has been the most applied one to evaluate the dielectric constant in liquid water because most classical MD and AIMD simulations adopt Ewald summation technique in computing the electrostatic potential energies. In Ewald summation, the applied electric field on a system is constrained to zero by the vanishing electric potential drop across the macroscopic sample, which is equivalent to applying the metallic boundary condition. Under the metallic boundary condition, the dipole fluctuation of water is transverse in nature. Therefore, a converged simulation of transverse dipole fluctuation in water is essential for an accurate prediction of its static dielectric constant. As shown in Fig. 2 (a), we present the electric dipole $\mathbf{M}$ as a function of simulation time along one of the orthogonal directions in the DPMD simulations. Clearly, it can be seen in Fig. 2 (a) that $\mathbf{M}$ has a mean value $\mathbf{M}=0$, which is expected since water is a paraelectric material with vanishing net polarization at room temperature. Nevertheless, $\mathbf{M}$ undergoes a large fluctuation and a significant second moment $\left\langle\mathbf{M}^{2}\right\rangle$ giving rise to its large static dielectric constant. However, $\left\langle\mathbf{M}^{2}\right\rangle$ converges rather slowly with simulation time as shown in Fig. 2 (c), which is due to the slow relaxation of polarization. According to Debye’s theory of dielectric relaxation, the polarization fluctuation decays exponentially over time whose characteristic time $\tau^{\epsilon^{\prime}=\infty}_{\mathbf{M}}$ can be determined from the autocorrelation $C_{\mathbf{M}}\left(t\right)$ of the electric dipole in the system defined as $C_{\mathbf{M}}\left(t\right)=\frac{\left\langle\mathbf{M}\left(t\right)\cdot\mathbf{M}\left(0\right)\right\rangle}{\left\langle\mathbf{M}\left(0\right)^{2}\right\rangle}$. In our simulation, a relaxation time $\tau^{\epsilon^{\prime}=\infty}_{\mathbf{M}}=8.1$ psps has been extracted from our DPMD simulation, which is close to the experimentally reported Debye relaxation time of 8.28 $ps$ at room temperature [40]. In the MD computer simulations, the simulation time $t$ should be long enough to sample sufficient cycles of polarization relaxation in water for a well-converged dipole fluctuation $\left\langle\mathbf{M}^{2}\right\rangle$. The formulas of the relative error $I_{\varepsilon-1}$ in the predicted dielectric constant are shown in Eq. 3 in which the accuracy is inversely proportional to $\frac{t}{\tau_{\mathbf{M}}}$ as expected. Because of the slow Debye relaxation, it has been widely noticed that simulation time at the scale of tens of nanoseconds is required to have a converged prediction of the static dielectric constant of water. In our DPMD simulation over 20 $ns$ , the dielectric constant of water is determined to be $\varepsilon=102$ with a small relative error of $3\%$ according to Eq. 3 [22].
@@ III.2 Dipole fluctuation under insulating boundary condition
The insulating electric boundary condition is equivalent to an open-circuit capacitor setup in experiments as schematically shown in Fig. 1 (b), in which no free charge exists on the metal plates. Because of the vanishing free charge on the capacitor, a constrained zero electric displacement field is imposed on the macroscopic sample of water [41]. Under the insulating boundary condition, the dipole fluctuation is longitudinal in nature. In comparison with the transverse dipole fluctuation under metallic boundary condition, the longitudinal dipole fluctuation is largely suppressed as shown in Fig. 2 (a). The mechanism is similar to that behind the experimentally observed higher phonon energies of longitudinal optical phonons compared to the transverse optical in ionic solid materials [42]. In addition to the suppressed dipole fluctuation, the polarization relaxation time $\tau^{\epsilon^{\prime}=0}_{\mathbf{M}}$ is greatly reduced as well. In a large band gap insulator such as neat water, no excess free charge $\rho_{\text{free}}$ exists, therefore $\nabla\cdot\mathbf{D}=\rho_{\text{free}}=0$ resulting in a uniform electric displacement field $\mathbf{D}$ throughout the system. The application of insulating electric boundary condition in DPMD simulation enables us to treat the long-range Coulombic interaction by the uniform parameter $\mathbf{D}$ throughout the neat water. Under this treatment, the dipole fluctuation in water becomes nearsighted and less correlated in real space, which results in a largely suppressed polarization relaxation time $\tau^{\epsilon^{\prime}=0}_{\mathbf{M}}$. As shown in Fig. 2 (b), the polarization relaxation time $\tau^{\epsilon^{\prime}=0}_{\mathbf{M}}=0.3$ ps is obtained in our computer simulations under the insulating electric boundary condition, and this value is in close agreement with the recent result reported by Sprik et al. [43]. Under insulating electric boundary condition, the fast polarization relaxation time $\tau^{\epsilon^{\prime}=0}_{\mathbf{M}}$ also helps the quicker convergence of $\left\langle\mathbf{M}^{2}\right\rangle$ as compared to that under metallic boundary conduction as seen in Fig. 2 (c). Because of the faster polarization relaxation time $\tau_{\mathbf{M}}^{\varepsilon^{\prime}}$ in insulating boundary condition, one expects that the predicted static dielectric constant of water should converge much faster than that in metallic boundary condition. However, our DPMD simulation suggests the opposite. As can be clearly seen in Fig. 2 (d), the predicted dielectric at the time scale of 20 $ns$ is barely converged with significant error bars. This seemingly surprising result is due to the fact that the prediction of $\varepsilon$ under insulating boundary condition demands the inversion of a small numerical number in Eq. 2 as $\varepsilon=\frac{1}{1-\frac{4\pi\left\langle\mathbf{M}^{2}\right\rangle}{3Vk_{\mathrm{B}}T}}$ which poses an additional challenge on the accuracy of predicted $\varepsilon$. The above effect can be more easily seen in the analytic formula of the relative errors in Eq. 2. The relative error of the insulating boundary condition has an extra term of $\varepsilon$ reflecting the loss of accuracy by inverting a small number. Based on Eq. 3, it can be estimated that a DMPD trajectory over 2.2 microseconds ($\mu s$) is required in order to achieve a similar accuracy as reported for a DPMD simulation performed under the metallic boundary condition within 20 ns.
@@ IV Dipole Correlation Method
In the above, the dipolar pair correlation function $c_{m}^{\varepsilon^{\prime}}\left(r\right)$, as defined in Eq. 6, is an ensemble function of distance $r$ normalized by the molecule number density $\rho$ and total number of molecules $N$ in the simulation cell, which accounts for the angular correlation between the unit vector of two dipole moments $\hat{\bm{\mu}}_{i}$ and $\hat{\bm{\mu}}_{j}$ on two distinct water molecules located at distances of $\bm{r}_{i}$ and $\bm{r}_{j}$ from the reference molecule at the origin point, respectively.
@@ IV.2 Long-Range Dipole Correlations in the H-Bond Network
The behavior of the dipolar correlation function $c_{m}^{\varepsilon^{\prime}}\left(r\rightarrow\infty\right)$ is a direct consequence of the long-range Coulomb forces. Therefore, an explicit treatment of the long-range electrostatic forces in the underlying deep neural network is essential to predict the correct asymptotic behavior. As shown in Fig. 5 (a) and (b), we compare the dipolar correlation functions $c_{m}^{\varepsilon^{\prime}=\infty}\left(r\right)$ generated by molecular simulations using the long-range Coulomb interaction inclusive DPLR model and the standard short-range DPSR model under the metallic boundary condition. In both DPLR and DPSR based molecular dynamics simulations, the liquid water structure is modeled by various sizes of simulation boxes of lengths $L$ that contain $N$ number of water molecules. Clearly, it can be seen that the DPLR model correctly predicts the constant of $c_{m}^{\varepsilon^{\prime}=\infty}\left(r\right)$ as $r\rightarrow\infty$. Moreover, the value of $c_{m}^{\varepsilon^{\prime}=\infty}\left(r\rightarrow\infty\right)$ decays with the volume of the simulation box, which scales as $\sim 1/V=1/L^{3}$ as expected, shown in the insert of Fig. 5 (a). In sharp contrast, $c_{m}^{\varepsilon^{\prime}=\infty}\left(r\rightarrow\infty\right)$ predicted by the DPSR model is qualitatively incorrect, yielding an exponential decay as $r\rightarrow\infty$ for all simulations, as shown in Fig. 5 (b). Indeed, in the absence of long-range Coulombic forces, the dipole correlation function, as an electronic property, exhibits exponential decay with distance from a local perturbation, in accordance with Walter Kohn’s nearsightedness principle [44].